OpenStudy (anonymous):

A line with slope 1/5 passes through (6,3). If the ordinates of a point on the line is 4, what is the abcissa? A line with slope 1/5 passes through (6,3). If the ordinates of a point on the line is 4, what is the abcissa? @Mathematics

OpenStudy (anonymous):

did you mean slope =0.2, trough point 6,3 and the other y is 4? the answer must be x=11, isnt it?

OpenStudy (anonymous):

did u get that/

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

slope= y2-y1/x2-x1 if my answer is not misunderstanding wit your question. 0.2=(4-3)/x-6 x=11